# A Survey of Secure Semantic Communications

### v1

@@ 5.4 Differential Privacy-based Deep JSCC
While differential privacy offers robust mathematical foundations and the capability to quantify and analyze privacy leakage risks, ensuring the accuracy and efficiency of SemCom while safeguarding privacy presents a formidable challenge. The incorporation of differential privacy may introduce additional data noise, potentially compromising the precision and efficiency of SemCom. Moreover, as datasets expand in size and queries become increasingly intricate, computational complexity escalates, further elevating the requirements for implementing differential privacy in SemCom systems. Consequently, future research endeavors will concentrate on integrating differential privacy techniques into deep JSCC processes, aiming to effectively extract semantic information while preserving privacy. Additionally, to mitigate the potential performance degradation associated with differential privacy in SemCom, future research will delve into optimization strategies for privacy protection mechanisms, encompassing adaptive differential privacy and local differential privacy.
@@ 5.5 Efficient Homomorphic Encrypted SemCom
Although homomorphic encryption offers the potential to achieve privacy protection without significantly impacting the performance of SemCom, its high complexity can pose challenges for certain terminals. Future research endeavors will focus on optimizing encryption algorithms and minimizing computational overhead. This includes exploring novel encryption algorithms that support a broader range of operations and fine-tuning encryption parameters, such as key length and ciphertext size, to enhance execution efficiency. Additionally, efforts will be made to reduce computational overhead by leveraging multi-core processors or distributed computing systems, as well as designing specialized hardware, such as FPGAs and ASICs, to expedite execution.
@@ 5.6 Smart Contract-Enabled SemCom
Research on SemCom empowered by smart contracts encompasses a range of areas, including data access control, verification of SemCom protocols, and anomaly detection. Specifically, this involves the design of permissions and rules for data access, ensuring that only users or devices meeting specific conditions can access sensitive semantic data. Furthermore, smart contract rules will be devised for verifying SemCom protocols, aiming to enhance the reliability and efficiency of the communication process. Additionally, strategies for monitoring and analyzing SemCom data will be developed to detect anomaly patterns, such as unusual semantic data traffic, transmission delays, or signs of semantic data tampering. Upon detecting such anomalies, smart contracts can swiftly trigger alerts and implement corresponding security measures, such as isolating affected nodes or terminating SemCom sessions to safeguard the integrity and security of SemCom systems.
@@ 5.7 Semantic Channel Fingerprint Database-Enabled PLS
The semantic channel fingerprints in the semantic channel fingerprint database can serve as a basis for authentication. When a device attempts to access the network, it only needs to compare its semantic channel characteristics with the information in the database to quickly determine its legitimacy. This significantly improves the efficiency and speed of identity authentication. Furthermore, the semantic channel characteristics in the semantic channel fingerprint database can be used as a basis for key generation. By extracting and analyzing the semantic channel characteristics of the device, related keys can be generated, thereby enabling secure key generation and distribution. In the future, the construction of the semantic channel fingerprint database will involve the extraction, description, and optimization of key semantic feature points, as well as the rapid retrieval, matching, and update management of the semantic channel fingerprint database.
@@ 7 Acknowledgements
The work presented in this paper was supported by the National Key R&D Program of China No. 2020YFB1806900, the National Natural Science Foundation of China No. 61871045, No. 62401074, and No. 61932005, Beijing Natural Science Foundation No. L242012 and No. L232051, the research foundation of Ministry of Education-China Mobile under Grant MCM20180101, and the Joint Research Fund for Beijing Natural Science Foundation and Haidian Original Innovation under Grant L232001.

### v2

@@ 5.4 Differential Privacy-based Deep JSCC
While differential privacy offers robust mathematical foundations and the capability to quantify and analyze privacy leakage risks, ensuring the accuracy and efficiency of SemCom while safeguarding privacy presents a formidable challenge. The detailed discussions are as follows.
Research Challenges:(1) Balancing privacy and accuracy: Differential privacy effectively safeguards user privacy, yet the accompanying data noise may undermine the accuracy and efficiency of SemCom. SemCom systems hinge on precise data interpretation and processing. Compromised data quality can lead to erroneous extraction of semantic information, thereby impacting the overall system performance. (2) Elevated computational complexity: As datasets grow larger and queries become more intricate, the computational load on SemCom systems intensifies. In this context, implementing differential privacy can significantly escalate computational resource consumption, potentially slowing the system’s response and detracting from user experience. (3) Necessity for privacy protection mechanism optimization: Differential privacy technology in SemCom is still nascent. Optimizing these mechanisms to mitigate their impact on system performance is a crucial area for future exploration. Examples include refining adaptive and local differential privacy technologies to better align with diverse data attributes and system demands.
Potential Solutions: (1) Optimizing differential privacy algorithms: To minimize data noise and ensure semantic information accuracy, differential privacy algorithms must be refined. Techniques like Gaussian Differential Privacy offer robust privacy protection without compromising accuracy. Furthermore, research should focus on enhancing the parallelization and distributed processing capabilities of these algorithms to handle large-scale data more efficiently. (2) Adaptive differential privacy: Developing adaptive differential privacy algorithms that dynamically adjust protection levels based on data characteristics and query requirements is essential. In high-risk situations, protection should be heightened; in low-risk contexts, it can be appropriately lowered to balance privacy and system performance. Real-time monitoring of data sensitivity and usage scenarios allows for dynamic adjustment of privacy protection parameters, informed by users’ historical behaviors and current requests. (3) Local differential privacy: Embracing local differential privacy technology distributes privacy protection across various nodes or devices. This reduces reliance on central servers and enhances system fault tolerance and scalability. In SemCom systems, nodes collaborate, mutually monitor, and protect privacy, fostering a multi-layered privacy protection mechanism.
@@ 5.5 Efficient Homomorphic Encrypted SemCom
Homomorphic encryption offers privacy protection capabilities for SemCom, facilitating direct computation on encrypted data to safeguard against sensitive information leakage. Nevertheless, its practical deployment is hindered by high computational complexity and substantial resource demands.
Research Challenges: The complexity of ciphertext operations in homomorphic encryption, including addition and multiplication, surpasses that of plaintext operations by a wide margin, leading to considerable delays in semantic encoding and decoding, and thus posing a challenge to meeting real-time communication requirements. For example, a single multiplication operation in FHE can take several seconds. Mobile terminals and IoT devices, constrained by limited computing power and memory, struggle to handle the intensive computational tasks associated with homomorphic encryption, thereby impeding the universal applicability of SemCom.
Potential Solutions: (1) Algorithm optimization and lightweight design: Dynamically adjust the encryption level according to the complexity of semantic tasks. For instance, employ FHE solely for key semantic features while opting for lightweight encryption for other components. Leverage semantic sparsity to devise sparse ciphertext computation protocols, thereby omitting encryption operations in non-critical areas. (2) Hardware-algorithm co-acceleration: Leverage FPGA/ASIC acceleration and design hardware that incorporates instruction sets tailored for homomorphic encryption, thereby hardware-accelerating core operations. Furthermore, harness the parallel computing prowess of GPUs to expedite large-scale ciphertext matrix operations.
@@ 5.6 Smart Contract-Enabled SemCom
Blockchain smart contracts, by leveraging their decentralized nature, transparent execution, and tamper-proof characteristics, offer data credibility assurance, automated protocol management, and robust anomaly response capabilities for SemCom. Nevertheless, their seamless integration with SemCom is hindered by various challenges as follows.
Research Challenges: (1) Execution efficiency and real-time constraints of smart contracts: The blockchain consensus mechanism and contract execution necessitate multi-node verification, which increases end-to-end latency in SemCom and poses difficulties in meeting real-time interaction demands, such as those in autonomous driving or industrial control. (2) Accuracy and efficiency challenges in anomaly detection: Complex anomaly detection algorithms cannot be directly implemented within smart contracts, and relying on off-chain computing might introduce trust risks. Furthermore, rule-based simple anomaly detection methods (like threshold determination) lack the capability to adequately identify attacks.
Potential Solutions: (1) Hierarchical architecture and on-chain/off-chain collaboration: Deploy core logic on-chain while offloading complex semantic processing tasks to a trusted off-chain execution environment. For instance, in IoV, real-time semantic interactions between vehicles are facilitated through state channels, with hash values of collision warning results stored on-chain for evidentiary purposes. (2) Efficient anomaly detection and response framework: Develop a lightweight rule base utilizing a finite state machine to identify common semantic attack patterns. AI anomaly detection models should be run on edge nodes, with only summaries of suspicious events submitted to on-chain contracts for verification.
@@ 5.7 Semantic Channel Fingerprint Database-Enabled PLS
The semantic channel fingerprints in the semantic channel fingerprint database can serve as a basis for authentication. When a device attempts to access the network, it only needs to compare its semantic channel characteristics with the information in the database to quickly determine its legitimacy. This significantly improves the efficiency and speed of identity authentication.
Research Challenges: (1) Environmental dynamic interference affecting channel fingerprints: The wireless channel fingerprints undergo substantial alterations due to multipath effects and moving obstructions. In IIoT, the movement of robotic arms leads to significant fluctuations in CSI, posing a challenge for traditional static channel fingerprint databases to manage effectively. (2) Risks of tampering in semantic channel fingerprint repositories: Attackers have the potential to inject forged fingerprints by generating CSI perturbations using GANs, which can impair the semantic extraction capability of channel fingerprints and mislead the authentication results.
Potential Solutions: (1) Implement spatial-temporal joint modeling: Achieve intelligent integration of dynamic CSI matrices and semantic features. (2) Enhance PLS with cryptographic techniques: Leverage blockchain-based distributed fingerprint evidence storage, utilizing an advanced Practical Byzantine Fault Tolerance (PBFT) consensus mechanism combined with Pedersen commitments to conceal the original features.
@@ 7 Acknowledgements
The work presented in this paper was supported by the National Key R&D Program of China No. 2020YFB1806905, the National Natural Science Foundation of China No. 61871045, No. 62401074, and No. 61932005, Beijing Natural Science Foundation No. L242012 and No. L232051, the research foundation of Ministry of Education-China Mobile under Grant MCM20180101, and the Joint Research Fund for Beijing Natural Science Foundation and Haidian Original Innovation under Grant L232001.